OpenStudy (anonymous):

I need help... solve 4a/9=100/45 a)5 b)25 c)-5 d)-25

OpenStudy (anonymous):

its A

OpenStudy (anonymous):

4a/9 = 100/45 Multiply both sides by 9 4a = 900 / 45 Work out right side 4a = 20 Divide both sides by 4 a = 5

OpenStudy (anonymous):

5 is your answer.

OpenStudy (anonymous):

thank yall sooo much

OpenStudy (anonymous):

You're welcome